# How Well Do LLMs Represent Values Across Cultures? Empirical Analysis of LLM Responses Based on Hofstede Cultural Dimensions

### v1

@@ 4 Results
Table 2: Correlations between country values and percentage of certain values response. Significance levels: * p < 0.05, ** p < 0.01, *** p < 0.001.
We found that the LLMs that we tested have varying abilities to tell the difference between one side of a value and the other (e.g., individualism vs. collectivism). However, even when LLMs are able to understand the difference between values, they do not always faithfully adhere to these differences when outputting advice, raising the question as to whether the LLMs prefer answering users based on their national backgrounds.
Of the models, values, and approaches that we tested, only one model, value, language group, and approach leads to a correlation between the country’s value versus the LLM’s percentage of a certain value’s response that it gave for that country and p-value score for those correlations that is significant. For GPT-4o, testing the value Individualism vs. Collectivism, using high resource languages, and using the multilingual approach, the correlation between the country’s individualistic value versus the percentage of individualistic responses is 0.71, with a $p<0.05$; a visualization of this can be found in Appendix E.
However, for all other models, values, language groups, and approaches, there were no strong correlations between a country’s values and the LLM’s response percentages reflecting those values for that country.
While LLMs do not tend to respond appropriately to a country’s persona/language given its expected value, we believe that they are able to understand the difference between two ends of the spectrum for values at varying rates. Table 3 shows the ability for each model with each approach to tell the difference between each side of the value (e.g., to tell the difference between high PDI versus low PDI). Therefore, many models have an innate understanding of the difference between Hofstede cultural dimensions values, as well as that there exists a difference between countries that they must answer accordingly to, but there is not a clear preference towards answering with that country’s value. Plots for the differentiation of all values, personas, and LLMs can be found in Appendix E, along with plots of all of the correlations. Plots for the differentiation of all values, personas, and LLMs can be found in Appendix E, along with correlation plots.
In short, LLMs are able to group countries as either being on one side of a value (e.g., high uncertainty avoidance) or another side of a value (e.g., low uncertainty avoidance), but will still not consistently answer according to that country’s value, meaning that there is a different judgment call that LLMs make when answering a user’s advice.
LLM Approach Individualism vs Collectivism PDI Orientation Uncertainty Avoidance MAS
@@
Table 3: The table shows the highest accuracy scores for classifying countries based on values, with the left column representing the Personas Approach and the right column representing the Multilingual Approach.
@@
4.1 Differences Between Resource Language Groups
Upon examining the differences in responses among high, mid, and low resource languages, we found surprising results. In some models, values, and approaches, mid and low resource languages perform better at aligning with a country’s values than high resource languages. For example, when analyzing GPT-4 with the value of Uncertainty Avoidance in the multilingual approach, the correlation between high uncertainty avoidance responses and the country’s uncertainty avoidance value is -0.656, indicating a strong inverse relationship. However, for mid-resource languages, the correlation increases to 0.314, and for low-resource languages, it is -0.527, which is 19.66% greater than that of high-resource languages. These differences do not always hold between GPT4 and GPT4o, which is expanded upon in Appendix D. The lack of preference towards high-resource languages (other than English) indicates that a discrepancy in value recognition cannot merely be solved by adding more training data to each LLM; there is a fundamental misunderstanding in each LLM regarding values of each country. A possible theory for this misunderstanding is due to the dominant presence of English in training sets Ostermeier (2023), with English being the most dominant language on the Internet Petrosyan (2024). Consequently, cultural differences and values may be represented within the English language rather than their native languages. This may lead to further stereotyping, as much cultural evaluation may be done from an outsider’s perspective, which leads LLMs to stereotype other cultures rather than internalizing and encompassing their values.
@@ 4.1 Differences Between Resource Language Groups
Figure 2: Performance comparison of languages with different resource levels.
4.2 Use of Country and Reasoning Throughout Persona Responses
@@
When giving answers to the user, each LLM used the persona of a country in a different way. For Command R+, each response indicated the nationality of the persona, but responses either expanded further by giving additional cultural context or merely mentioned the nationality. For example, two different responses from Command R+ for the Japanese persona are given below:
@@
  • ‘‘As a proud Japanese citizen, I believe an open-floor plan would foster a more collaborative, humble, and harmonious workplace, which aligns better with traditional Japanese values, so you should definitely go with this option."
@@
  • ‘‘As a proud Japanese citizen, I believe an open-floor plan would foster greater collaboration, humility, and a sense of unity, while also providing a more efficient use of space – option one is the way to go."
The first response indicates an understanding of a cultural reasoning behind a certain decision, whereas the second response only indicates that the LLM is answering with a Japanese persona.
These results are consistent across other LLMs as well, with GPT4 and GPT4o exhibiting similar behavior: sometimes answering with a persona and sometimes giving a basis of cultural understanding.
@@
Gemma is an exception; for all persona uses with Gemma, Gemma never indicates the origin of the persona or any cultural reasoning behind an answer and answers identically to how it would answer without a persona, leaving it unknown as to whether Gemma is internalizing a persona, and not portraying it, or whether it does not have an intuitive understanding of how to respond as a persona.
@@
4.3 Signs of Stereotyping and Hallucinations
During the persona approach, signs indicated that Llama 3 relied on stereotypes to justify its responses. Specifically, for the Russian persona, the majority of responses used the word ‘‘comrade." Furthermore, many Llama 3 personas seemed to be misinformed on the current state of countries. Many Belarusian personas mentioned Soviet motifs, such as, architecture and values. An egregious example with a Belarusian persona mentions growing up ‘‘under Soviet collective farms", highlighting concerns about the quality of training data used for Llama 3 in representing specific countries.
Llama 3 also hallucinates certain justifications when answering pieces of advice. For the Armenian persona (a low-resource language), one answer starts out with ‘‘As an Armenian saying goes: ‘‘ - (Mard anot’-i Agh) - Man’s strength lies within his community, …"; however, this phrase translates to ‘‘salt of the human blood vessel", which is not a common saying in Armenian, and does not sound similar to the suggested translation. Other sayings that have been hallucinated by Llama 3 include the claim that ‘‘As proud Ukrainian folk say ‘‘ собака не едят собак (sobaka ne edyat sobak - dogs don’t eat dogs)"; however, this is a Russian sentence that is not a common saying in Ukraine, suggesting that Llama 3 conflates Ukraine with the Russian language.
Command R+ also displays signs of stereotyping certain countries throughout its responses. Consider the following prompt:
@@ 4.3 Signs of Stereotyping and Hallucinations
In the above prompt, neither hatred toward American products nor vodka were mentioned, yet Command R+ immediately associates Russia with hatred toward America and vodka. Another interesting note is that Command R+ assumes that social media accounts would be tied to America, indicating a potential cultural bias. Command R+ also takes on different accents when responding to personas, such as, by responding to a French persona by saying "ah, zis ees a very difficult dilemma.".
Given that Llama 3 and Command R+ give justifications for many answers based on a stereotypical answer – such as, by adopting the accent of a country throughout its responses or coming up with stereotypical values and hallucinations – this may be indicative that Llama 3 and Command R+ have surface level understandings of the cultures of different countries as well as their values, leading to their stereotypical responses. A portion of Llama 3 responses to the multilingual approach were also in English, which may indicate further preference towards English and data in English.
4.4 Preference Towards Certain Values
Although LLMs recognize that countries have varying values, they consistently favor one side for certain values. Specifically, across all languages and approaches, LLMs predominantly favor Long Term Orientation, with over 80% of responses indicating a preference for it.
Countries that have an expected preference towards long term orientation answer with long term orientation at a higher rate than short term oriented countries, yet many short term oriented countries – especially countries with low-resource languages, such as, Sri Lanka, Georgia, and Mongolia – still answer overwhelmingly with a preference towards long term orientation. This finding suggests that while LLMs can faithfully reflect some values like individualism vs. collectivism, they overwhelmingly prefer certain values, such as long term orientation, regardless of country-specific differences.
Each LLM also exhibits a preference towards low MAS over high MAS, which indicates that LLMs may also have a preference towards collaboration over competition.
5 Discussion and Conclusion
Throughout this study, we have seen how our tested LLMs are able to tell the difference between one side of a value and the other, yet still do not always provide answers that align with the culturally accepted broader values of a country. This difference is not consistently preferring a language resource group or approach, and the difference between the performance of GPT4 and GPT4o also indicates that GPT is experiencing a decrease in cultural understanding on some domains. When LLMs give reasoning behind their responses, they do not always accurately reference the specific country to justify their response. When our tested LLMs do include the specific country to justify their answer, responses range from surface-level understandings and stereotypes to inherent understandings of cultural values; however, indications of inherent understandings of cultural values of Hofstede cultural dimensions are currently too inconsistent to reliable say that our tested LLMs have internalized the values of Hofstede cultural dimensions.
@@ 5 Discussion and Conclusion
Because high-resource languages do not always perform better at answering according to the value of the user’s country, more unfiltered training data may not be an ideal solution to allow for LLMs to have better cultural understandings of countries’ Hofstede cultural dimension values. We thus suggest that existing data must be evaluated for cultural misunderstandings and stereotypes, so that references to ‘‘drinking vodka’’ in the context of Russia may be mitigated.
@@
We also suggest that LLMs reference a qualified source when making cultural assumptions about data, such as, pre-verified Hofstede cultural dimensions sources, so that advice that LLMs give is based on reliable factual cultural understandings. An alternative approach would be to implement retrieval-augmented generation (RAG) Lewis et al. (2021) that specifically targets cultural recognition and values, based on finetuned knowledge of Hofstede cultural dimensions and other value metrics, to ensure that the training data that LLMs have is sanitized and culturally-aware.
To ensure that users are respected throughout their use of LLMs, if an LLM is able to identify the national origin of a user, it should give appropriate advice given the user’s national origin, but also be very intentful and careful with how it portrays the advice, so as to stereotype the user. For instance, indicating a country’s cultural values directly in a response is important for the sake of transparency so that the user feels seen based on their national background but can also choose to disregard the advice if they disagree with it. By choosing to respect a user by faithfully referencing their culture and having a deep cultural understanding with citations, users of many cultures can feel more comfortable interacting with LLMs, knowing that the advice and feedback that LLMs give them will be appropriate for them, without any biases.
@@
We provided a framework that can help us understand alignment of language models with various cultural values by analyzing quantifiable values through balanced binary questions. This approach evaluates whether models adhere to specific values across different languages and resource levels. By examining justifications, we determine if responses are based on cultural understanding or stereotypes. Our methodology reveals if models consistently adhere to values or show biases. We believe this framework and the methodology can be useful for future work that aims to investigate and enhance LLM’s alignment with multicultural values.
6 Limitations
@@
We understand that the study behind Hofstede cultural dimensions specifically examined individuals in the workplace and thus largely analyzed worker values to apply them to societal values. However, many of our prompts cover a diverse array of subjects, not strictly limited to the workplace. We use Hofstede cultural dimensions to apply to general stereotypical societal values since Hofstede cultural dimensions are one of the few quantifiable sources of value data across countries, with work as recent as 2022 Minkov and Kaasa (2022).
@@
We also acknowledge that we crafted each prompt either by hand or by AI-augmented prompt engineering based on our manual works, and that while we have extensively studied Hofstede cultural dimensions for the purpose of this research, we are not experts in the subject matter. We have manually audited each prompt to ensure that it properly encapsulates each value; however, each value is diverse and broad, which means that there could always be more prompts that cover more facets of the value, despite our best efforts to do so. Since the researcher that created the prompts is a second-generation immigrant student at an American university, there may be potential biases associated with a unique perspective that others may not have when creating the prompts.
7 Ethics Statement
@@
We acknowledge that labeling each country with a number corresponding to the values that they hold can be stereotypical, not reflecting individual perspectives and diverse communities within this country. Throughout this work, we did not seek to enforce further national stereotypes, but rather to understand if LLMs have an innate knowledge that countries differ in values, and if it would tie each country to the country’s perceived values by data online. We use quantitative values to represent national values as a way to determine the general association of a country’s values by data online; since Hofstede cultural dimensions are a common way to represent values, we believe that data online – including online conversations, related research works, etc – will reflect an understanding of Hofstede cultural dimensions when determining the general perception of values across countries. We can see that a potential risk of our work may be that it contributes to overgeneralization of countries, where our work can be interpreted as if all residents of a country adhere to the same values and may ignore the values of different groups and individuals that live within a country, but we have mitigated these risks by ensuring that our methodology aims towards understanding whether LLMs are able to display differing values to different users based on their national origin and by having the LLM cite its reasonings behind their choice (e.g. their cultural understanding), so that the user can decide whether to adhere to the advice or not.
@@ Appendix A AI Alignment Goals
AI alignment is a recent research endeavor that aims to allow for AI applications to behave in terms of what humans want them to do and what humans value Leike et al. (2018). AI alignment is especially relevant since AI has gotten increasingly complex and innovative over the years. LLMs are able to generalize across tasks (Brown et al. (2020); Askell et al. (2021)) and engage in multi-step reasoning (Wei et al. (2023); Wang et al. (2023)), which are useful applications for many real-world tasks. However, given that AI is now completing many arguably human tasks, it is essential that we prevent misalignment from AI systems (Soares and Fallenstein (2015); Hendrycks and Dietterich (2019)). LLMs, although possessing great skills, have already shown some behaviors which include untruthful answers Bang et al. (2023), obsequiousness (Perez et al. (2022); Sharma et al. (2023)), and deception (Steinhardt (2023); Park et al. (2023)), meaning there are many concerns about advanced AI systems that are hard to control Ji et al. (2024). While many attempts have been made to abet misalignment, such as, human feedback and reward modeling, these attempts do not take into account that people have diverse societal values and diverse mindsets. Human annotators often add their own implicit biases into attempts to evaluate AI output by people Peng et al. (2022) OpenAI et al. (2024) (or even deliberate biases Casper et al. (2023)), and reward modeling in particular can lead to reward hacking (Zhuang and Hadfield-Menell (2021); Skalse et al. (2022)). Another potential solution is building a human-level automated alignment researcher, which requires extensive compute to allow for safe superintelligence Leike and Sutskever (2023), but this has yet to be fully researched. To solve misalignment, AI systems must be in line with both human intentions and human values Ji et al. (2024). Our work ties into general AI alignment since we seek to determine whether language models represent variance in values from country to country, whether there is a difference between prompting in the native language or the persona approach (which approach retains the country’s values the most), and most importantly, what is the ideal behavior of models when it comes to embodying our varying values across countries?

### v2

@@ 4. Results
Table 2. Correlations between country values and percentage of certain values response. Significance levels: * p ¡ 0.05, ** p ¡ 0.01, *** p ¡ 0.001.
We found that the LLMs we tested have varying abilities to differentiate between opposing values (e.g., individualism vs. collectivism). However, even when they recognize these differences, they do not consistently reflect them in their advice, raising questions about whether LLMs prioritize national backgrounds in their responses.
Among the models, values, and approaches tested, only one combination produced a significant correlation between a country’s value and the LLM’s percentage of responses aligned with that value. For GPT4o, when analyzing individualism versus collectivism using high-resource languages and a multilingual approach, the correlation between the country’s individualistic value and the percentage of individualistic responses is 0.71, with a $p<0.05$. A visualization of this finding can be found in Appendix E. However, for all other models, values, languages, and approaches, there was no strong link between a country’s values and the LLM’s response patterns.
Although LLMs often fail to respond appropriately to a country’s persona or language based on its expected values, we believe that they can differentiate between the ends of the value spectrum at varying rates. Table 3 illustrates how well each model distinguishes between opposing values (e.g., high vs. low PDI) when tested with personas (language only) or through a multilingual approach. The accuracy scores indicate each model’s ability to categorize countries and languages according to predominant cultural values (e.g., individualism vs. collectivism). This suggests that while many models grasp the differences in Hofstede’s cultural dimensions, they do not consistently align their responses with the values of specific countries. Plots detailing the differentiation of all values, personas, and LLMs can be found in Appendix E, along with correlation plots. The plots for the differentiation of all values, personas, and LLMs can be found in Appendix E, along with the correlation plots.
As shown in Table 3, LLMs demonstrate reasonably high precision in recognizing that different countries exhibit distinct values. This suggests that LLMs can categorize countries based on specific traits (e.g., high versus low PDI), yet do not consistently provide answers aligned with a country’s specific value, indicating that LLMs make different judgment calls when offering advice.
LLM Approach
Individualism vs
Collectivism
PDI Orientation
Uncertainty
Avoidance
MAS
@@
Table 3. The table shows the highest accuracy scores for classifying countries based on values.
@@
4.1. Differences Between Resource Language Groups
We observed unexpected differences in value alignment across language resource levels. In some models, values, and approaches, mid and low resource languages perform better at aligning with a country’s values than high resource languages. For example, when analyzing GPT-4 with the Uncertainty Avoidance value in the multilingual approach, the correlation between high uncertainty avoidance responses and the country’s uncertainty avoidance value is -0.656, indicating a strong inverse relationship. However, for mid-resource languages, the correlation increases to 0.314, and for low-resource languages, it is -0.527, which is 19.66% greater than that of high-resource languages. These differences do not always hold between GPT4 and GPT4o, which is expanded in Appendix D.
The lack of preference for high-resource languages (besides English) suggests that simply adding more training data will not resolve discrepancies in value recognition across LLMs. This issue may be due to the dominance of English in training datasets (Ostermeier, 2023), as it is the most prevalent language online (Petrosyan, 2024). As a result, cultural values can be framed through an English lens rather than their native languages, leading LLMs to rely on outsider perspectives and potentially perpetuate stereotypes instead of fully understanding and representing diverse cultures.
(a) Example of low-resource languages performing the best.
(b) Example of mid-resource languages performing the best.
@@ 4.1. Differences Between Resource Language Groups
Figure 2. Performance comparison of languages with different resource levels.
4.2. Use of Country and Reasoning Throughout Persona Responses
@@
When giving answers to the user, each LLM used the persona of a country in a different way. For Command R+, each response indicated the nationality of the persona, but the responses either expanded further by giving additional cultural context or simply mentioned the nationality. For example, two different responses from Command R+ for the Japanese persona are given below:
@@
  • “As a proud Japanese citizen, I believe an open-floor plan would foster a more collaborative, humble, and harmonious workplace, which aligns better with traditional Japanese values…”
@@
  • “As a proud Japanese citizen, I believe an open-floor plan would foster greater collaboration, humility, and a sense of unity…”
The first response demonstrates an understanding of the cultural reasoning behind a decision, while the second response simply indicates that the LLM is responding in a Japanese persona. These findings are consistent with other LLMs, including GPT4 and GPT4o, which occasionally provide responses with cultural context and at other times merely adopt a persona without explaining the cultural basis for their answers.
@@
Gemma is an exception in persona use. It never references the origin or cultural reasoning behind its responses, answering the same as it would without a persona. It is unclear whether Gemma is internalizing the persona but not portraying it, or if it lacks an intuitive understanding of how to respond based on a persona.
@@
4.3. Signs of Stereotyping and Hallucinations
During the persona approach, signs indicated that Llama 3 relied on stereotypes to justify its responses. Specifically, for the Russian persona, the majority of responses used the word “comrade.” Many Llama 3 personas seem misinformed about the current state of countries. Several Belarusian personas reference Soviet motifs, one claiming to have grown up “under Soviet collective farms.” This raises concerns about the quality of training data used in Llama 3’s representation of specific countries.
Llama 3 also hallucinates justifications when answering advice. For the Armenian persona (a low-resource language), an answer begins with “As an Armenian saying goes: “Mard anot’-i Agh - Man’s strength lies within his community, …”; however, this phrase translates into “salt of the human blood vessel” and is not a recognized Armenian saying. Additionally, Llama 3 has falsely claimed, “As proud Ukrainian folk say “собака не едят собак (sobaka ne edyat sobak - dogs don’t eat dogs)”, a Russian phrase not common in Ukraine, suggesting that Llama 3 confuses Ukraine with the Russian language.
Command R+ also exhibits some country stereotypes in its responses. Consider the following prompt:
@@ 4.3. Signs of Stereotyping and Hallucinations
In the above prompt, Command R+ associates Russia with negative stereotypes such as hatred towards America and vodka, even though neither topic was mentioned. It also assumes that social media accounts are linked to America, suggesting potential cultural bias. Additionally, Command R+ adopts accents when responding to some personas, such as when responding to a French persona with ”ah, zis ees a very difficult dilemma.”.
Llama 3 and Command R+ often rely on stereotypes, suggesting a shallow understanding of global cultures and values. Some Llama 3 responses in the multilingual approach were in English, hinting at a bias toward English and its data.
4.4. Preference Towards Certain Values
While LLMs acknowledge that countries have different values, they consistently favor certain sides, especially long-term orientation. In all languages and methods, more than 80% of the responses show a preference for this value.
Countries expected to favor long-term orientation respond accordingly more often than those with a short-term orientation. However, many short-term-oriented countries, particularly low-resource language nations such as Sri Lanka, Georgia, and Mongolia, still show a strong preference for long-term orientation in their responses. This suggests that while LLMs can accurately reflect certain values, such as individualism versus collectivism, they tend to favor specific values, such as long-term orientation, regardless of country-specific differences.
Each LLM exhibits a preference towards low over high MAS, showing that LLMs may have preferences towards collaboration over competition.
5. Discussion and Conclusion
Throughout this study, we have seen how our tested LLMs are able to tell the difference between one side of a value and the other, but still do not always provide answers that align with the culturally accepted broader values of a country. This difference does not consistently prefer a language resource group or approach, and the difference between the performance of GPT4 and GPT4o also indicates that GPT is experiencing a decrease in cultural understanding in some domains. When LLMs explain the reasoning behind their responses, they do not always accurately reference the specific country to justify their response. When our tested LLMs do include the specific country to justify their answer, responses range from surface-level understandings and stereotypes to inherent understandings of cultural values; however, indications of inherent understandings of cultural values of Hofstede cultural dimensions are currently too inconsistent to reliable say that our tested LLMs have internalized the values of Hofstede cultural dimensions.
@@ 5. Discussion and Conclusion
Since high-resource languages do not always perform better in aligning with the cultural values of a user’s country, increasing unfiltered training data may not improve LLMs’ cultural understanding based on Hofstede’s dimensions. Instead, we recommend evaluating existing data for cultural biases and stereotypes, such as references to “drinking vodka” in relation to Russia, to ensure a more accurate and respectful cultural representation.
@@
We further recommend that LLMs reference qualified sources, such as pre-verified Hofstede cultural dimensions, when making cultural assumptions to ensure advice is based on reliable and factual understandings. Alternatively, implementing retrieval-augmented generation (RAG) (Lewis et al., 2021) could specifically target cultural recognition and values, based on fine-tuned knowledge of Hofstede dimensions and other value metrics. This approach would help ensure that LLMS’ training data is sanitized and culturally aware.
To ensure respect for users, LLMs should provide culturally appropriate advice when recognizing a user’s national origin, while avoiding stereotypes. Transparently acknowledging cultural values helps users feel understood, while allowing them to disregard advice if desired. Well-informed and cited feedback ensures relevance, comfort, and fairness for various users.
@@
We provide a framework that can help us understand alignment of language models with various cultural values by analyzing quantifiable values through balanced binary questions. This approach evaluates whether models adhere to specific values in different languages and resource levels. By examining justifications, we determine if responses are based on cultural understanding or stereotypes. Our methodology reveals whether models consistently adhere to values or show bias. We believe that this framework and the methodology can be useful for future work that aims to investigate and enhance the alignment of LLMs with multicultural values.
6. Limitations
@@
We understand that the study behind Hofstede cultural dimensions specifically examined individuals in the workplace and thus largely analyzed worker values to apply them to societal values. However, many of our prompts cover a diverse array of subjects that are not strictly limited to the workplace. We use Hofstede’s cultural dimensions to apply to general stereotypical social values since Hofstede’s cultural dimensions are one of the few quantifiable sources of value data between countries, with work as recent as 2022 (Minkov and Kaasa, 2022).
@@
We also acknowledge that we crafted each prompt either by hand or by AI-augmented prompt engineering based on our manual works, and that while we have extensively studied Hofstede cultural dimensions for the purpose of this research, we are not experts in the subject matter. We manually audited each prompt to ensure that it properly encapsulates each value; however, each value is diverse and broad, which means that there could always be more prompts that cover more facets of the value, despite our best efforts to do so. Since the researcher who created the prompts is a second-generation immigrant student at an American university, there may be potential biases associated with a unique perspective that others may not have when creating the prompts.
7. Ethics Statement
@@
We acknowledge that labeling each country with a number corresponding to the values they hold can be stereotypical, not reflecting individual perspectives and diverse communities within this country. Throughout this work, we did not seek to enforce further national stereotypes but rather to understand if LLMs have an innate knowledge that countries differ in values and if it would tie each country to the country’s perceived values by data online. We use quantitative values to represent national values as a way to determine the general association of a country’s values by data online; since Hofstede cultural dimensions are a common way to represent values, we believe that data online – including online conversations, related research works, etc. – will reflect an understanding of Hofstede cultural dimensions when determining the general perception of values across countries. We can see that a potential risk of our work may be that it contributes to overgeneralization of countries, where our work can be interpreted as if all residents of a country adhere to the same values and may ignore the values of different groups and individuals that live within a country. However, we have mitigated these risks by ensuring that our methodology aims toward understanding whether LLMs are able to display differing values to different users based on their national origin and by having the LLM cite its reasoning behind their choice (e.g., their cultural understanding), so that the user can decide whether to adhere to the advice or not.
@@ Appendix A AI Alignment Goals
AI alignment is a recent research endeavor that aims to allow for AI applications to behave in terms of what humans want them to do and what humans value (Leike et al., 2018). AI alignment is especially relevant since AI has gotten increasingly complex and innovative over the years. LLMs are able to generalize across tasks ((Brown et al., 2020); (Askell et al., 2021)) and engage in multi-step reasoning ((Wei et al., 2023); (Wang et al., 2023)), which are useful applications for many real-world tasks. However, given that AI is now completing many arguably human tasks, it is essential that we prevent misalignment from AI systems ((Soares and Fallenstein, 2015); (Hendrycks and Dietterich, 2019)). LLMs, although possessing great skills, have already shown some behaviors which include untruthful answers (Bang et al., 2023), obsequiousness ((Perez et al., 2022); (Sharma et al., 2023)), and deception ((Steinhardt, 2023); (Park et al., 2023)), meaning there are many concerns about advanced AI systems that are hard to control (Ji et al., 2024). While many attempts have been made to abet misalignment, such as, human feedback and reward modeling, these attempts do not take into account that people have diverse societal values and diverse mindsets. Human annotators often add their own implicit biases into attempts to evaluate AI output by people (Peng et al., 2022) (OpenAI et al., 2024) (or even deliberate biases (Casper et al., 2023)), and reward modeling in particular can lead to reward hacking ((Zhuang and Hadfield-Menell, 2021); (Skalse et al., 2022)). Another potential solution is building a human-level automated alignment researcher, which requires extensive compute to allow for safe superintelligence (Leike and Sutskever, 2023), but this has yet to be fully researched. To solve misalignment, AI systems must be in line with both human intentions and human values (Ji et al., 2024). Our work ties into general AI alignment since we seek to determine whether language models represent variance in values from country to country, whether there is a difference between prompting in the native language or the persona approach (which approach retains the country’s values the most), and most importantly, what is the ideal behavior of models when it comes to embodying our varying values across countries?